OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4 2 3 All real numbers No solution

OpenStudy (anonymous):

No solution

OpenStudy (anonymous):

WHAT ABOUT THIS ONE Solve for x: 5x + 2 = 4x − 9 7 −7 11 −11

OpenStudy (s):

-2(x+3)=-2(x+1)-4 -2x-6=-2x-2-4 -2x+2x=6-2-4 0x=6-6 0x=0 so there is no solution

OpenStudy (s):

the next one : 5x + 2 = 4x − 9 5x-4x=-9-2 x=-11

OpenStudy (anonymous):

can u help me with this one One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single? 576 240 144 960